OpenStudy (anonymous):

(103.5) (32.5 - 24.5) (4.181) = (31.5) (102.4 - 32.5) (x) solve for X

OpenStudy (callisto):

(103.5) (32.5 - 24.5) (4.181) = (31.5) (102.4 - 32.5) (x) (103.5) (8) (4.181) = (31.5) (69.9) (x) x = [ (103.5) (8) (4.181) ] / [ (31.5) (69.9) ] = Can you work it out?

OpenStudy (anonymous):

\[\Large (103.5)(32.5-24.5)(4.181)=(31.5)(102.4-32.5)(x)\] \[\Large \frac{(103.5)(32.5-24.5)(4.181)}{(31.5)(102.4-32.5)}=x\]

OpenStudy (callisto):

Honestly, what you need to do now is to take out your calculator, press the numbers and get the answer!